OpenStudy (anonymous):

find the equations of the two tangent lines of the graph of f that pass through the indicated points. f(x)=x^2 with points on it at (-1,-1),(3,) with an outside point at (1,-3)

OpenStudy (anonymous):

(3,9) sorry left out the 9

OpenStudy (anonymous):

ok you want the line tangent to the graph of \[y=x^2\] at the point \[(3,9)\]? there is only one, so not sure what that other condition is supposed to mean

OpenStudy (anonymous):

are you sure this is right? (3,9) is on the graph because \[f(3)=3^2=9\] so there is only one tangent line

OpenStudy (anonymous):

on the other hand \[(-1,-1)\] is not on the graph, so you can find a line through (-1,-1) tangent to the graph. that is a different problem all together

OpenStudy (anonymous):

i have this graphed out (x^2) and it has two points on it which are the ones i put above. with tangent lines going through them and they intersect at the (1,-3) point

OpenStudy (anonymous):

oh and i meant (-1, 1)

OpenStudy (anonymous):

ok \[f(-1)=1\] so that point is on the graph of \[y=x^2\] the tangent line at that point has slope -2, since the derivative of x^2 is 2x the line will therefore have the equation \[y-1=-2(x+1)\] or \[y=-2x-1\] but there is only one line tangent to the curve there.

OpenStudy (anonymous):

likewise at (3,9) the slope will be 6, and so the line will be \[y-9=6(x-3)\]

OpenStudy (anonymous):

how do you know the slope is 6?